Subject: 3.5-m Users Committee minutes 5/05/08
From: Bruce Gillespie
Submitted: Tue, 6 May 2008 08:03:03 -0600
Message number: 1102
(previous: 1101,
next: 1103
up: Index)
APO 3.5-m Users Committee Phonecon, 5/05/08
Attending: Suzanne Hawley, Bruce Gillespie, Scott Anderson, John
Bally, Mark Klaene, Jon Holtzman, Michael Strauss, Russet McMillan,
and Remy Indebetouw
Absent: Al Harper, Jon Fulbright
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Strauss) - Michael said the PU users were happy.
UVa (Indebetouw) - Remy said that there was "no unhappiness" among UVa
users.
UW (Anderson) - Scott mentioned that two UW users have recently had
trouble starting TUI on Red Hat (4) platforms. The apparent workaround
is to log off and then log back on. Russell Owen is looking into the
problem.
CU (Bally) - John said he had heard nothing back from the CU users.
JHU (Fulbright) - no report.
NMSU (Holtzman) - Jon had nothing new to report on user feedback or
issues.
UChicago (Harper) - no report.
**********************************
Discussion of telescope/instruments report - Klaene:
The report is given below. Mark provided updates on the forest
closure and other items. There have been two fires burning both east
and north of APO, although they are not likely to threaten the
observatory. Recent high winds continue to degrade utility electrical
power, and the backup generator, although usable, needs servicing
sometime in the next couple of months. The telescope has been
generally well-behaved, with some minor problems addressed in the
report. Although the autofocus script is being tested, other work
priorities will delay its initial release to users for a few months.
We have recently adopted a modified dust policy: the Observing
Specialists will open the telescope under borderline high dust
conditions, and then assess the dust levels inside the open enclosure
(with the pressurization fans running to help keep the air clean near
the telescope) and use the in-enclosure dust readings to determine
whether to stay open or not. This has enabled us to recently keep the
telescope open under conditions that we would have normally not opened
the telescope in the past, based solely on outside dust levels.
The replacement of the fan-out board for NIC-FPS is nearly completed;
all indications are positive so far. The final instrument cool-down
was delayed because the ion pump appeared to be broken, and it was
replaced with a spare. Once NIC-FPS is cooled down, tests of the noise
performance of the detector will be made, possibly in the next day or
two. The N-F Fabry-Perot etalon was removed, and will be returned to
CU for cryo-testing. If the etalon is still inoperative in the lab, it
will be returned to the vendor for repairs. Parts needed for moving
Agile to a permanent port (with a rotator) were received and are being
assembled. We should be ready to move Agile in a couple of months.
Related to the recent problem with dust, Bruce mentioned that Mike
Skrutskie had asked if we have any data on the K-band emissivity of
the telescope (we don't). Now that we have two low-noise NIR
instruments, telescope emissivity related to temperature and
cleanliness of the optics is especially worthy of our attention. Mike
suggested that we might be able to baseline and monitor telescope
emissivity using NIC-FPS and/or TripleSpec. Everyone agreed that this
is a good idea that someone (who?) should pursue.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 3/24/08
through 4/29/08
Mark Klaene & Bruce Gillespie
0) Overview
Operations continued to be largely routine through the period. The
spring weather (with its seasonal high winds and elevated dust levels)
has affected observing--we lost approximately 4.5 nights due the high
winds and dust levels during this period. The fire danger level in the
surrounding forest is now extremely high, and although the forest is
going to be "closed" effective May 1 by the USFS, no impact to normal
operations is expected.
On site-related matters, recent water usage at the site has been
unusually high, and two underground water leaks were subsequently
located and repaired. However, there appears to be a third leak
somewhere that requires further investigation. Site electric power has
been relatively unstable, due in part to the high winds that have
caused grid fluctuations and outages. Although operational, our
emergency backup generator has a fuel leak for which we are trying to
locate a part.
1) Telescope
Tertiary rotation has become problematic again. The engineering staff
is working hard to diagnose the problem, but progress is slow because
the symptoms do not appear during daytime testing. Slippage of the
main azimuth drive system was noticed during tracking early last week,
and the problem was fixed with a slight preload adjustment and the
servicing of the lower azimuth bearing. We started telescope and
enclosure preventive maintenance, those annual items that can be done
without a shutdown. An autofocus script is now available for initial
testing by the Observing Specialists. The script uses the NA2 guider
focus to adjust the secondary, and has had some limited testing to date.
2) Instruments
DIS, SPIcam, and the Echelle operated smoothly throughout the period.
Echelle experienced a minor warm-up that was traced to a loose power
connection. Shortly afterwards, the Echelle ion pump controller
failed. We believe this was the source of the burnt odor noticed two
months ago. A new ion pump controller was installed, and no further
warm-up problems have been seen. The new fan-out board replacement for
NIC-FPS and the etalon removal work has started.
With the arrival of TripleSpec, CorMASS has been decommissioned and
crated for shipment back to UVa in the upcoming weeks. The visiting
instruments (GFP and APOLLO) were both operational. TripleSpec
commissioning began and the new instrument is well on the way to being
available for scheduled science. Some TripleSpec first-light
information is available--contact John Wilson at UVa if you are
interested. We are also planning to place preliminary TripleSpec user
documentation on the web soon.
3) CIF projects
We are making good progress on the implementation of Agile field
rotation and its migration to the TR2 mirror cell port. Most of the
new parts have been received and initial assembly is underway, to be
followed by bench testing. The tertiary support/actuator upgrade is
progressing; long-lead items have been purchased, and machining has
started. Design is slightly behind schedule but still on track for an
August installation. The rotator drive upgrade has been in routine
operational use for some time now, and indications are that its
performance requirements appear to have been met. Ongoing cleanup of
rotator controller software will continue over the coming months,
which will enable the migration of the new axis controller system to
the telescope altitude drive.
4) Summer Shutdowns
There will be three shutdown periods this summer. Two 3-day shutdowns,
one in June and one in July, will be for secondary and axis drive
work. In August, we are planning a 17-day shutdown for the tertiary
support upgrade, as well as for M1 washing and several additional
maintenance tasks (e.g., annual instrument servicing).
**********************************
Q3 Summer shutown dates - Hawley:
Suzanne said that that the major summer shutdown will occur 6 - 22
August (inclusive), and a mini-shutdown will be scheduled for 14 - 16
July.
**********************************
TripleSpec update - Indebetouw, Hawley:
Remy said that UVa is pleased with the progress in commissioning
TripleSpec. The online user documentation should be posted this week,
and the data reduction software is very near completed and ready for
users. Suzanne added that TripleSpec had two engineering runs during
March and April, and that the TUI control functions appear to be in
place. The instrument will be offered to the general user community
this summer in a shared-risk, block-scheduled manner. The blocks will
be scheduled in 4 or 5 contiguous nights, so that UVa staff can be on
call to field questions. Data taken during the coming quarter will be
user-proprietary. Russet added that since we are block-scheduling
TripleSpec to enable UVa support, it will not be available as a fill-
in or back-up instrument outside of the scheduled blocks. If users
wish to write a TripleSpec proposal for the coming quarter, Suzanne
said that there is a link on the APO instrument information webpage to
the UVa documentation, and people should contact John Wilson (jcw6z at
virginia.edu) if there are questions not covered in the documentation.
In addition, TripleSpec data already taken during the engineering runs
will continue to be available for prospective users to look at;
contact Suzanne or Bruce for the URL. Suzanne asked that the User
Committee members should make sure that their users are aware that
TripleSpec is being made available in the coming quarter.
**********************************
NIC-FPS update - Bally:
In addition to news about NIC-FPS that is covered in the reports
above, Stephane Beland and Mike Kaiser (from CU, CASA) have been at
APO installing the new fan-out board and removing the F-P etalon. In
the next day or two, full-up tests of the detector will be possible to
see if the new wiring improves the detector noise characteristics.
**********************************
ACTION ITEMS:
[open from previous months]:
ACTION: John Bally will write a section with figures on Fowler
sampling for inclusion in the NIC-FPS user documentation.
STATUS: Closed. John recently reviewed the section and made editorial
comments, which were implemented by Gabrelle Saurage. In general,
Suzanne asked that we remind all users that we depend on their
feedback for improving the 3.5-m telescope, instrumentation, and
software user documentation on an ongoing basis.
[new actions from this month]:
none
**********************************
Future of the Photometric Telescope - Hawley, Gillespie:
Although not on the agenda, Suzanne and Bruce reported on an item
discussed at the recent ARC Board of Governors telecon. The Board
agreed with the plan to assume operational responsibility of the PT
after SDSS-II ends on 14 July. Proposals for post-SDSS programs for
the PT have been received and are being evaluated. At the November
annual Board meeting, we will make a recommendation on which future
program(s) to conduct on the PT, with a plan for managing and funding
said program(s).
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 9 Jun, at 8:30 AM
Pacific Time. The agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1102 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO